OpenStudy (123ab456c):

Northside Picnic Grounds is shaped like a triangle. Two perpendicular sides of the grounds measure 1.8 miles and 2.5 miles. What is the area of the picnic grounds? A. 2.15 mi² B. 2.25 mi² C. 4.3 mi² D. 4.5 mi²

OpenStudy (123ab456c):

can you tell me the answer and then explain it that way i will understand

OpenStudy (welshfella):



OpenStudy (123ab456c):

why do you type so slowwww

OpenStudy (welshfella):

the 2 sides marked are perpendicular so that the base is 1.8 and the height is 2.5 and the area of a triangle = 1/2 * base * height which in this case is 1/2 * 1.8 * 2.5

OpenStudy (123ab456c):

is the answer D

OpenStudy (welshfella):

No try again 1/2 * 1.8 * 2.5 = 0.9 * 2.5 = ?

OpenStudy (123ab456c):

OHHH the answer is B

OpenStudy (welshfella):

yes looks like you didn't multiply by 1/2 the first time.

OpenStudy (123ab456c):

yeah! tiny mistake thanks!

OpenStudy (welshfella):

yw